OpenStudy (anonymous):

what is 7/5 times 4/25 times 5/14 times 3/4 as a simplified fraction the answer i got, mathlab says its wrong

OpenStudy (anonymous):

You should have \[\frac{3}{50}\]

OpenStudy (curry):

3/50

OpenStudy (anonymous):

\[\frac{7}{5}\frac{4}{25}\frac{5}{14}\frac{3}{4}\] The 4s and 5s cancel. The 7 and 14 cancel leaving 2 where the 14 was and 1 where the 7 is: \[\frac{1}{1}\frac{1}{25}\frac{1}{2}\frac{3}{1}=\frac{3}{50}\]

OpenStudy (anonymous):

my math work is online and it says 3/50 is incorrect

OpenStudy (anonymous):

Programs are picky. How is the question worded?

OpenStudy (anonymous):

7/5 times 4/25 times 5/14 times 3/4 as a simplified fraction

OpenStudy (anonymous):

It's possible it's a bugged question then. May want to raise that one with your instructor for adjustment.

OpenStudy (anonymous):

ok ill do that thanx for the help